a cubical box of 3m edge contains cubes having 30cm of edges.how many number of cubes does the box contain

Answer by mananth(16946)

volume of box = l^3= 3*3*3=27 m^3
edge of cube = 30 cm = 0.3 m
volume of cube = 0.3*0.3*0.3
=0.027 m^3
number of cubes= volume of box/ volume of each cube
27/0.027
=1000 cubes